sumitra has Rs 34 in 50 paise and 25 paise coins. If the number of 25 paise coins is twice the number of 50 paise coins , how many coins of each kind does she have.
Answer by mananth(16949)

50 paise coins = x
25 paise coins = 2x
She has Rs34 = 3400 paise
50*x+25(2x)=3400
50x+50x=3400
100x=3400
x=3400/100
x= 34 ---- 50 paise coins
25 paise coins ---- 34*2= 68
CHECK
68*25 + 34*50 = 3400
m.ananth@hotmail.ca